OpenStudy (anonymous):

Evaluate each function:

OpenStudy (anonymous):

\[h(t)=\left| t+2 \right| +3\] ; find h(6)

OpenStudy (anonymous):

Just change t to 6. \[h(6) = |6+2| + 3 = |8| + 3 = 8 + 3 = 11\]

OpenStudy (anonymous):

Haha(: thanx! Do you mind helping with 2 more?? If so: \[\prime(n)=n^2-2n\] Find \[\prime(n^2)\]

OpenStudy (anonymous):

You have to change n to n². So, you will got: \['(n^2)=(n^2)^2-2(n^2)=n^4-2n^2\]

OpenStudy (anonymous):

Thnx!! and last but not least: \[g(a)=4a\] ; find \[g(2a)\]

OpenStudy (anonymous):

Again, change a to 2a. \[g(2a)=4(2a)=8a\]

OpenStudy (anonymous):

mmkay(: I think I got the hang of it now(: Thanks!! GBU!!(: